# LatentEditor: Text Driven Local Editing of 3D Scenes

### v2

@@ 3.2.2 NeRF Editing
Figure 4: Qualitative Results. The visual results of our approach, when contrasted with the baseline IN2N [9] across two distinct scenes, distinctly demonstrate that LatentEditor excels in accurately pinpointing the pertinent region, executing faithful text edits, and averting undesired alterations. These achievements prove challenging for the baseline method [9] to replicate effectively as IN2N [9] also changes the background objects’ color to blue given that editing prompt, “Give the tyres blue color” only wants the tyre color to be changed.
Figure 6: Addressing Complex Prompt. Demonstrating the multi-attribute editing capability in 3D NeRF Scenes – Unveiling the versatility and precision of our approach through intricate modifications and complex prompt editing in the 3D environment.
@@ 4 Experimental Analysis
Figure 9: Qualitative comparison of edited image with and without refinement loss against the ediitng prompt, “turn his hair black”. The red box indicates the noticeable artifacts without the refinement module.
@@ 4.1 Results
Quantitative evaluations, as detailed in Table 2, involved 56 edits across 14 scenes from the above mentioned four datasets. Our method outperformed baselines in CLIP similarity scores and CLIP direction consistency, as averaged over multiple views rendered from NeRF. This superiority in scoring demonstrates the precision of LatentEditor in aligning edited NeRF scenes with target text conditions.
@@ F.2 NeRF Editing - Efficiency Analysis
Figure 3: Comparison between DreamEditor and LatentEditor. Our results are more realistic. This approach not only preserves the attributes of the original object but also maintains the background color intensity consistent with the original.
@@ G Additional Results
Figure 5: LatentEditor Qualitative results on LLFF dataset  (see the provided videos).
@@ G.1 Qualitative Results
We further present the qualitative comparisons with Control4D [41] and DreamEditor [54] in Figures 23, and 4. Figure 2 illustrates the contrasts in local and multi-attribute editing among LatentEditor, IN2N, and Control4D. As the Control4D [41] code is not publicly accessible, we developed an in-house implementation to assess the impact of GAN guidance. Control4D tends to alter the original scene content, while IN2N inadvertently changes unintended areas, like beard and jacket color when only a hair color change is intended. Conversely, LatentEditor precisely achieves the editing objectives as per the prompts, demonstrating its superior accuracy in targeted modifications.
In Figure 3, the comparison between DreamEditor and LatentEditor is shown. LatentEditor delivers more realistic outcomes and maintains the background color intensity consistent with the original scene, as can be seen in the provided videos. This fidelity in preserving the authenticity of the scene underscores the effectiveness of LatentEditor in nuanced editing.
Moreover, Figure 4 compares LatentEditor with NeRF-Art [48], DreamEditor [54], and IN2N [9], focusing on style transfer capabilities. LatentEditor distinguishes itself by achieving the desired style transfer results in significantly fewer training iterations, about 3000, compared to IN2N, which requires approximately 18000 iterations. This efficiency in obtaining high-quality outcomes with fewer iterations highlights the practical utility of LatentEditor in style transfer applications. Figure 5 shows the editing on a scene from the LLFF [28] dataset where we are editing the color of the television.
@@ G.3 Comparing with Mask+IP2P
We further evaluate LatentEditor by combining a pre-trained segmentation model SAM [16] with IN2N [9] and IP2P [2]. The quantitative results are presented in Table 3

### v3

@@ 3.2.2 NeRF Editing
Figure 4: Qualitative Results. The visual results of our approach, when contrasted with the baseline IN2N [9] across two distinct scenes, distinctly demonstrate that LatentEditor excels in accurately pinpointing the pertinent region, executing faithful text edits, and averting undesired alterations. These achievements prove challenging for the baseline method [9] to replicate effectively as IN2N [9] also changes the background objects’ color to blue given that the editing prompt, “Give the tyres blue color” only wants the tyre color to be changed.
@@
Figure 8: Qualitative comparison of an edited image with and without refinement loss against the editing prompt, “turn his hair black”. The red box indicates the noticeable artifacts without the refinement module.
@@ 4.1 Results
Quantitative evaluations, as detailed in Table 2, involved 56 edits across 14 scenes from the above-mentioned four datasets. Our method outperformed baselines in CLIP similarity scores and CLIP direction consistency, as averaged over multiple views rendered from NeRF. This superiority in scoring demonstrates the precision of LatentEditor in aligning edited NeRF scenes with target text conditions.
@@ G.1 Qualitative Results
We further present the qualitative comparisons with Control4D [41] and DreamEditor [54] in Figure 2. Figure 2 illustrates the contrasts in local and multi-attribute editing among LatentEditor, IN2N, and Control4D. As the Control4D [41] code is not publicly accessible, we developed an in-house implementation to assess the impact of GAN guidance. Control4D tends to alter the original scene content, while IN2N inadvertently changes unintended areas, like beard and jacket color when only a hair color change is intended. Conversely, LatentEditor precisely achieves the editing objectives as per the prompts, demonstrating its superior accuracy in targeted modifications.
Moreover, Figure 3 compares LatentEditor with NeRF-Art [48], DreamEditor [54], and IN2N [9], focusing on style transfer capabilities. LatentEditor distinguishes itself by achieving the desired style transfer results in significantly fewer training iterations, about 3000, compared to IN2N, which requires approximately 18000 iterations. This efficiency in obtaining high-quality outcomes with fewer iterations highlights the practical utility of LatentEditor in style transfer applications.
@@ G.3 Comparing with Mask+IP2P
We further evaluate LatentEditor by combining a pre-trained segmentation model SAM [16] with IN2N [9] and IP2P [2]. The quantitative results are presented in Table 3.